this is my problem
x+3y=-4
2x+2y=-1
and my question says solve by the elimination methos.. whats the answer?

Answer by scott8148(6628)

multiplying the 1st equation by -2 gives ___ -2x - 6y = 8

adding to the 2nd equation gives ___ -4y = 7 (x is ELIMINATED)

y = -7/4

substituting ___ x + 3(-7/4) = -4 ___ x = 5/4